OpenStudy (anonymous):

check answer are the power input and output in watts. If an amplifier has a power gain of 16 watts and the output power was 12 watts, find the power input in watts. answer 2.133

OpenStudy (anonymous):

16=1o log P(o)/P(i) P(o)/P(i)=10^1.6 p(i)=12/((!0)^1.6)=.3